The mean is u=60 and its stardard devation is q=4.0
Find the probability that X is less that 53

Answer by stanbon(75887)

The mean is u=60 and its standard deviation is q=4.0
Find the probability that X is less that 53
--
z(53) = (53-60)/4 = -7/4
---
P(x < 53) = P(z < -7/4) = normalcdf(-1000,-7/4) = 0.0401
============================================================
Cheers,
Stan H.